OpenStudy (anonymous):

Given f(x) = the quantity of 2x plus 4 divided by 5 solve for f-1(3). 1 5 7 8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

x = 3 2(3) -2/ 4 6 - 2 = 4 4/4 = 1?

OpenStudy (anonymous):

f(x) = \[2x - 2 \over 4\]

ganeshie8 (ganeshie8):

check ur f(x) again, in the wording u said : f 2x plus 4 divided by 5 but in the equaiton u have a minus

OpenStudy (anonymous):

its minus

OpenStudy (anonymous):

it like converted my equation into words for some reason... the equation i posted is the correct one

ganeshie8 (ganeshie8):

okie, they want u find f^-1(3)

ganeshie8 (ganeshie8):

not f(3) get that difference first

ganeshie8 (ganeshie8):

f^-1(3) and f(3) are two different things

OpenStudy (anonymous):

yeah I want f^-1(3)

ganeshie8 (ganeshie8):

f^-1(3) means for wat value of x, the function f(x) becomes 3

ganeshie8 (ganeshie8):

\(\large f(x) = \frac{2x - 2}{4}\)

ganeshie8 (ganeshie8):

\(\large 3 = \frac{2x - 2}{4}\)

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

so.. would I add 2?

OpenStudy (anonymous):

im a little confused lol

ganeshie8 (ganeshie8):

first multiply 4

ganeshie8 (ganeshie8):

\(\large 3 = \frac{2x - 2}{4}\) mutiply 4 both sides \(\large 12 = 2x - 2\)

ganeshie8 (ganeshie8):

rest shud be easy :)

OpenStudy (anonymous):

add 2? 14 = 2x /2 x = 7?

ganeshie8 (ganeshie8):

thats it ! remember below :- 1) f^-1 means they want u find x 2) f means they want u find f(x)

OpenStudy (anonymous):

alright thanks man :)